# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ 4.1 Launch of the kpc scale multiphase wind
Figures 6 and 7 provide detailed insights into the launch of the multiphase outflow in the starburst region of the FD simulation. Generally, the launch of galactic winds in our simulations occurs in two stages: the formation and breakout of a superbubble foam, followed by the development of a kpc-scale multiphase wind. Figure 8 presents an edge-on view of the second stage, showcasing the formation of the kpc-scale outflow.
The first stage, spanning from $t\sim 5$ Myr to $t\sim 15$ Myr, involves the formation and breakout of a superbubble foam. Initially, the turbulent ISM is dominated by cool gas residing in low-density voids, high-density filaments, and very dense clumps. A few isolated supernova remnants (SNRs) emerge around 5 Myr, expanding into the cool ISM. As shown in the top right panel of Figure 6, these SNRs form within dense clumps, leading to significant metal enrichment of the surrounding cool gas. The lower-density ISM in the voids is more easily ionized and expelled outward. As more SNe occur, their remnants merge to form a superbubble foam. Between $t=10$ and 12.5 Myr, a single superbubble dominates the central starburst region, filled with warm and hot gas and permeated by high-density cool filaments (see the temperature panels in Figures 6 and 7). This superbubble expands and eventually breaks through the dense gas disk before $t\sim 15$ Myr, with outflowing gas reaching heights of 500 pc.
Figure 6: Face-on view of the central region of simulation FD. The left, middle, and right columns show the gas density, temperature, and metallicity. From the top to the bottom row, results at time t=10, 12.5, 15.0, 17.5, and 20.0 Myr are shown.
Figure 7: Edge-on view of the central region of simulation FD. The left, middle, and right columns show the gas density, temperature, and metallicity. From the top to the bottom row, results at time t=10, 12.5, 15.0, 17.5, and 20.0 Myr are shown.
@@
Following the superbubble breakout, hot gas moves outward along the minor axis of the disk at the highest velocity, followed by warm gas, and then cool gas in tenuous filaments. Figure 7 illustrates the significant interaction and mixing between cool filaments and the hot phase in the outflow, which contributes to the acceleration of the cool phase (Schneider et al. 2020; Fielding & Bryan 2022). As the filaments are stretched, their lengths increase, reaching hundreds of parsecs by $t=15$ Myr. These filament features in the wind are remarkably similar to recent JWST observations (Bolatto et al. 2024).
While the visual appearance of these cool filaments in the wind bears some resemblance to previous simulations of M82’s wind (e.g., Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020, the formation and evolution of these filaments in our simulations differ from previous work, which often attributed their formation to the interaction between spherical cold clouds and the hot wind. For instance, Cooper et al. (2008) shows that small, spherical cold clouds, formed from the disruption of larger gas clouds, are entrained into the outflow. These spherical clouds are then stretched by ram pressure to form filaments.
@@
Our picture align with previous simulations of giant molecular cloud disruption at resolutions of 0.01-0.1 pc (see, e.g., Kim et al. 2018; Li et al. 2019; Fukushima & Yajima 2021). These studies demonstrate that feedback from massive stars ionizes and expels low-density gas within GMCs, while high-density filamentary structures can survive. However, our simulation’s 4 pc resolution limits its ability to resolve cloud structures on parsec scales. Therefore, both cool gas in filamentary structures with length of a few to 100 pc, and smaller, nearly spherical clouds may survive the feedback from the starburst. As the warm and hot phases expand outward, both cool filamentous and cloud-like structures can be entrained into the outflow, contributing to the formation of the observed cool filaments in the M82’s wind.
@@
The multiphase wind comprises not only hot and cool phases but also a considerable amount of warm gas. Some of this warm gas originates from the disk, while other portions arise from the mixing layer between the hot and cool phases. Additionally, stellar feedback, primarily from core-collapse SNe, effectively enriches the ISM in the central region. By $t=15$ Myr, a substantial fraction of the multiphase gas in the sub-kpc wind has been enriched to metallicities exceeding solar. Subsequently, even more gas in the wind attains super-solar metallicities
@@
After the super bubble foam breaks out, the second stage, the formation of a kpc-scale multiphase wind, takes another 10-15 Myr. From t = 15 Myr onward, the wind continues to expand outward. This multiphase outflow consists of volume-filling hot gas, mass-dominant cold gas, and warm gas. As shown in Figure 8, the multiphase outflow extends beyond 1.5 kpc by t = 20 Myr. By t = 25 Myr, the outflow in the fiducial run FD reaches the upper boundary of the simulation volume (2 kpc along the minor axis), with numerous cool filaments stretching beyond 1 kpc. The outflow base has a radius of approximately 1.5 kpc. While some gas may escape the simulation volume by t = 30 Myr, the limited volume size and simulation time restrict our ability to fully capture the long-term evolution of the outflow.
@@
Figure 9: The outward velocity distribution of cool ($\rm{T}<2\times 10^{4}\,\rm{K}$, top), warm ($2\times 10^{4}\,\rm{K}<\rm{T}<2\times 10^{5}\,\rm{K}$ middle) and hot ($\rm{T}>2\times 10^{5}\,\rm{K}$, bottom) gas at $t=30Myr$ at different distance, r, from the minor axis. From the left to the right columns, results at a height of 500, 1000, 1500, and 2000 pc are shown. The color indicates the mass outflow rate in each r-$v_{z}$ bin.
@@
Figure 10: Top: Blue lines indicate the net inflow rate of cool ($T<2\times 10^{4}\,\rm{K}$) gas into three zones of a height between 500 to 1000 pc (left), 1000 to 1500 pc (middle) and 1500 to 2000 pc as functions of time, while the yellow line indicate the time variations of cool gas mass in corresponding zone. Bottom: Same as the top row, but for the integrated net inflow rate and the total cool gas mass.
@@ 4.2 Evolution of different phases in the wind
Figure 9 shows the distribution of the velocity along the minor axis, denoted as $V_{z}$, for the cool ($\rm{T}<2\times 10^{4}\,\rm{K}$, top panel), warm ($2\times 10^{4}$ to $2\times 10^{5}$ K, middle panel), and hot ($\rm{T}>2\times 10^{5}$ K, bottom panel) gas phases at different distances from the minor axis, r, in the FD simulation at 30 Myr. From left to right, the panels show results at heights of 500, 1000, 1500, and 2000 pc. A significant difference in the $V_{z}$-r space from z=500 to z=2000 pc is observed between the cool gas and the other two phases. At $z=500$ pc, most of the cool gas has a vertical velocity between $\sim 80\,\rm{km\cdot s^{-1}}$ and  $\sim\,300\rm{km\cdot s^{-1}}$, which increases moderately with height. This velocity is well below the escape velocity, suggesting that most of the cool gas in the wind will eventually fall back onto the disk.
In comparison, Martini et al. (2018) found that the line-of-sight velocity of most HI in the M82 wind is around $150\,\rm{km\,s^{-1}}$ at $\rm{z}=0.5$ kpc and $50\,\rm{km\,s^{-1}}$ at $\rm{z}=1\,\rm{kpc}$. Considering that M82 has an inclination angle of $77\pm 3^{\circ}$, these correspond to vertical velocities $V_{z}$ around $550\,\rm{km\,s^{-1}}$ and $200\sim 300\,\rm{km\,s^{-1}}$, respectively. Observations of H-$\alpha$ emission indicate that the three-dimensional radial outflow velocity of gas with a temperature of $10^{4}$ K is roughly 150 km/s at $r\sim 0.5$ kpc, rising to 500-600 km/s at $r\sim 1.0$ kpc, and remaining relatively constant in the outer regions (Xu et al. 2023b). Our simulated velocities for the cool phase are approximately $30\%-50\%$ of the observed values.
@@
In contrast, the warm and hot gas in our simulations exhibits higher vertical velocities. Most warm gases have velocities between $300\sim 600\,\rm{km\cdot s^{-1}}$, while hot gas velocities range from $600\sim 1000\,\rm{km\cdot s^{-1}}$. A significant fraction of the warm and hot ionized gas in the wind can escape the galaxy’s potential well. However, the velocities of the hot gas in our simulations are still about $20\%-30\%$ lower than those inferred from X-ray observations of M82 (Strickland & Heckman 2009).
To investigate the origin of the cool gas in the outflow, we compare the net inflow rate with the time variation of the gas mass within three height zones along the minor axis: 500-1000 pc, 1000-1500 pc and 1500-2000 pc. Figure 10 shows that the net inflow rates closely match the growth rates of cool gas in these zones, indicating that most of the cool gas in the wind originates from the disk rather than cooling from the warm and hot phases. While some warm and hot gas can cool and contribute to the cool phase, as shown in the lower right panel of Figure 10, their overall contribution to the cool phase in the outflow is relatively minor.
@@
We also analyzed the metallicity distribution of the three phases in the outflow (figure not shown). The three phases exhibit similar metallicity distributions, with most gas having a metallicity of $Z\sim 0.5-2.0\,Z_{\odot}$, approximately 50-100 times higher than the initial metallicity in the FD simulation. This indicates that all gas in the outflow has undergone significant metal enrichment.
@@ 4.3 Overall feature of the outflow in various simulations
The launch of galactic winds varies significantly across different simulations. To investigate the impact of different feedback mechanisms, Figures 11 and 12 compare the face-on and edge-on views of the wind and disk at t = 25 Myr for simulations with different feedback configurations.
Compared to the FD simulation, the wind in the fSN simulation exhibits a slightly larger opening angle and higher gas density due to the increased total SN energy input. However, the intense SN feedback in the fSN simulation results in a more dispersed gas distribution in the nuclear region. In contrast, the nSN simulation fails to launch a galaxy-scale outflow. Figure 11 shows that the gas disk in the nSN simulation is less turbulent and more concentrated in multiple mini-arms compared to simulations with SN feedback. Although some smaller bubbles filled with warm gas, driven by stellar winds and radiation feedback, are present, only a small amount of gas can escape into the halo, reaching heights of a few hundred to 1 kpc (Figure 12).
When both stellar wind and radiation feedback are excluded (nFB simulation), the gas disk becomes even more compact. In this case, almost no gas from the disk can escape into the halo. Additionally, without SNe, the gas retains its initial metallicity. Figure 13 presents edge-on views of the outflow at 30 Myr for all simulations. Reducing the gas return fraction leads to a decrease in the mass of the outflow. A higher initial gas metallicity results in a larger outflow volume due to enhanced star formation and SN feedback. A less massive gas disk has a modest impact on the strength of the outflow.
@@
Figure 14: Left: the mass density function at $t=30$ Myr calculated along the minor axis in various simulations. Right: outflow rate at different heights along the minor axis.

### v4

@@ 4.1 Launch of the kpc scale multiphase wind
Figure 6: Face-on view of the central region of simulation FD. The left, middle, and right columns show the gas density, metallicity, and temperature. From the top to the bottom row, results at time t=10, 12.5, 15.0, 17.5, and 20.0 Myr are shown.
Figure 7: Edge-on view of the central region of simulation FD. The first, second, and third (from left to right) columns show the gas density, metallicity, and temperature, while the last column is the zoomed-in view of the temperature map overlapped with the velocity field
. From the top to the bottom row, results at time t=10, 12.5, 15.0, 17.5, and 20.0 Myr are shown.
@@
Generally, the launch of galactic winds in our simulations occurs in two stages: the formation and breakout of a superbubble foam, followed by the development of a kpc-scale multiphase wind. For the first stage, Figures 6 and 7 provide detailed insights into the launch of the multiphase outflow in the central starburst region of the FD simulation. As for the second stage, Figure 8 presents an edge-on view of the whole simulation zone, showcasing the formation of the kpc-scale outflow.
The first stage, spanning from $t\sim 5$ Myr to $t\sim 15$ Myr, involves the formation and breakout of a superbubble foam. Initially, the turbulent ISM is dominated by cool gas residing in lower-density ($n\lesssim 5\ \rm{cm}^{-3}$) voids, intermediate to high density filaments, and very dense giant clouds with size of tens pc to 100 pc. A few isolated supernova remnants (SNRs) emerge around 5 Myr, expanding into the cool ISM. These SNRs form within dense giant clouds, leading to significant metal enrichment of the surrounding cool gas, as shown by the panels in the second and third columns of Figure 6. The lower-density ISM in the voids is more easily ionized and expelled outward. As more SNe occur, their remnants merge to form a superbubble foam. In contrast, gas in intermediate to high density filaments and dense giant clouds are less affected by SNe feedback. A considerable fraction of the filaments can survive from the feedback at $t<15$ Myr, as shown by the density plots in Figures 6 and 7. Between $t=10$ and 12.5 Myr, a single superbubble is formed and dominates the central starburst region, filled with warm and hot gas, numerous intermediate to high density cool filaments, and a few cool dense giant clouds (see the density and temperature panels in Figures 6 and 7). This superbubble expands and eventually breaks through the dense gas disk before $t\sim 15$ Myr, with outflowing gas reaching heights of 500 pc.
Figure 9: Filaments and cores of the neutral gas ($\rm{T}<8000\rm{K}$) in a slice of thickness 200 pc between $t=20$ and $30$ Myr in the FD simulation. The plug-in plot in each panel presents a zoom-in view of the interface between disc and outflow region with a size of $\times750$ pc.
Figure 10: Similar to Figure 9, but for $\rm{H\alpha}$ emitting gas($8000\rm{K}<\rm{T}<23000\rm{K}$).
Figure 11: Shown is the overlapped image of the neutral filaments(red) and $\rm{H\alpha}$ emitting filaments(blue), both identified by eigenvalues of the Hessian matrix of gas density, at 30 Myr
Following the superbubble breakout, hot gas moves outward along the minor axis of the disk at the highest velocity, followed by warm gas and, subsequently, cool gas, which is primarily carried in numerous filaments. The density and temperature panels in Figures 6 and 7, especially the fourth column of Figure 7, which provides a zoomed-in view of the interface between the central starburst region in the disc and the outflow within a $250\times 250$ pc box, demonstrate that a significant fraction of cool gas in the disk region ($z<125$ pc), largely contained in intermediate to high density filaments, can survive early feedback ($t<10$ Myr) and subsequently be entrained into the wind at $t>12.5$ Myr. The zoom-in panels at $t\gtrsim 12.5$ Myr in Figure 7 capture a considerable number of cool filaments still residing in the disk ($z<125$ pc) as they begin to move outward.
@@
Meanwhile, the zoom-in velocity field reveals substantial differences between the hot, warm, and cool phases flowing out of the disk, which can trigger mixing and Kelvin–Helmholtz instabilities. In this environment, interactions between the hot and cool phases can facilitate significant momentum transfer, accelerating the cool gas in filaments as it moves outward (Gronke and Oh, 2018; Schneider et al., 2020; Fielding and Bryan, 2022). During this process, the cool filaments can further stretched by the hot wind, their lengths increase, reaching hundreds of parsecs by $t=15$ Myr. These filament features in the wind are remarkably similar to recent JWST observations (Bolatto et al. 2024).
The visual appearance of these cool filaments in the outflow resembles previous simulations of M82-like galaxies (e.g., Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020). However, in our simulation, the dominant net contributor to the cool filaments in the outflow is the initially filamentary cool ISM in the disc. This contrasts with several previous studies, which typically attributed filament formation to interactions between initially non-filamentary, cold, clumpy clouds and the hot wind. For instance, simulations by Cooper et al. (2008) show that small, cold clouds, produced by the disruption of larger gas clouds, are entrained into the outflow by hot wind. The entrainment and subsequent mixing with the hot wind both accelerates these initially non-filamentary clouds through momentum transfer and stretch them into filamentary structures (Gronke and Oh 2018; Fielding and Bryan 2022).
@@
Our picture aligns better with previous simulations of giant molecular cloud disruption at resolutions of 0.01-0.1 pc (see, e.g., Kim et al. 2018; Li et al. 2019; Fukushima and Yajima 2021). These studies demonstrate that feedback from massive stars ionizes and expels low-density gas within GMCs, while intermediate and high-density filamentary structures can survive. Though our simulation’s resolution is coarser than these studies, limiting the ability to resolve finer structures on parsec scales, similar behavior can be captured in our simulation: the pre-existing intermediate and high-density filamentary structures in the ISM can survive the early stage of SNe feedback. At later times, they are entrained into the outflow and gradually accelerated through momentum transfer during mixing with the hot wind, all while largely retaining their filamentary morphology (Gronke and Oh 2018; Mandelker et al. 2020; Bustard and Gronke 2022; Fielding and Bryan 2022).
@@
The multiphase wind comprises not only hot and cool phases but also a considerable amount of warm gas. Some of the warm gas likely originates from the disk, while the rest may form within the mixing layers between the hot and cool phases. Additionally, stellar feedback, primarily from core-collapse SNe, effectively enriches the ISM in the central region. By $t=15$ Myr, a substantial fraction of the multiphase gas in the sub-kpc wind has been enriched to metallicity exceeding solar. Subsequently, even more gas in the wind attains super-solar metallicity.
@@
After the superbubble foam breaks out, the second stage, the formation of a kpc-scale multiphase wind, takes another 10-15 Myr. From t = 15 Myr onward, the wind continues to expand outward. This multiphase outflow consists of volume-filling hot gas, mass-dominant cold and cool gas, and warm gas. As shown in Figure 8, the multiphase outflow extends beyond 1.5 kpc by t = 20 Myr. By t = 25 Myr, the outflow in the fiducial run FD reaches the upper boundary of the simulation volume (2 kpc along the minor axis), with numerous cool filaments stretching beyond 1 kpc. The outflow base has a radius of approximately 1.5 kpc. While some gas may escape the simulation volume by t = 30 Myr, the limited volume size and simulation time restrict our ability to fully capture the long-term evolution of the outflow.
To further explore the origin and evolution of the cool and warm phase in the outflow, particularly the formation and evolution of filamentary structures between $t=20$ to $30\rm Myr$, when these phases in the outflow develops most rapidly, we identify filaments using a method based on the eigenvalues of the Hessian matrix of the gas density field. This approach is commonly applied to filaments identification in both the cosmic web and the ISM (Hahn et al., 2007; Bond et al., 2010; Schisano et al., 2014; Zhu and Feng, 2017). In our analysis, grid cells with at least two positive eigenvalues are classified as belonging to filamentary (or core) structures and are further separated into two temperature components: neutral filaments ($T<8000\rm K$) and $\rm H\alpha$-emitting filaments ($8000\rm K<T<23000\rm K$). Figures 9 and 10 present the edge-on density field of the identified neutral filaments and $\rm H\alpha$-emitting filaments of the FD simulation from $t=20$ to $30\rm Myr$, both figures are averaged over $200\,\rm pc$-thick slice. Numerous filaments can be observed in both figures, while only a few dense cores appear.
Figures 9 and 10 show that most cool filaments originate in intermediate to high density regions of the gas disk (below $z<200$ pc), as highlighted in the zoom-in view of Figure 9. Numerous filaments, ranging from tens to several hundred parsecs in length, emerge from the disk, extend into the outflow region ($z>200$ pc), and assemble into web-like structures. Once launched into the wind, these filaments are partially accelerated and stretched by the surrounding high-velocity hot and warm gas. In this process, portions of the filaments can be reshaped into arc-like morphologies, reminiscent of the structures observed by JWST (Lopez et al., 2025). Figure 11 further illustrates the spatial correlations between neutral filaments, $\rm H\alpha$ filaments, and bubbles of high-velocity hot gas. Many of the $\rm H\alpha$ arcs are connected to neutral filaments, yet they often reside at the shock fronts of hot winds, suggesting that at least some $\rm H\alpha$ arcs originate from interactions between cool filaments and hot outflows.
@@
Figure 12: The outward velocity distribution of cool ($\rm{T}<2\times 10^{4}\,\rm{K}$, top), warm ($2\times 10^{4}\,\rm{K}<\rm{T}<2\times 10^{5}\,\rm{K}$ middle) and hot ($\rm{T}>2\times 10^{5}\,\rm{K}$, bottom) gas at $t=30Myr$ at different distance, r, from the minor axis. From the left to the right columns, results at a height of 500, 1000, 1500, and 2000 pc are shown. The color indicates the mass outflow rate in each r-$v_{z}$ bin. The solid and dashed orange lines denote the escape velocity calculated with and without DM halo, respectively. The red vertical line indicates the half mass radius of the gas disk.
Figure 13: Similar to Figure 12, shown is the metallicity distribution of the three phases at $t=30Myr$ at different position in the outflow region.
@@
Figure 14: Top: Blue lines show the net mass flux of cool gas ($T<2\times 10^{4}\,\rm K$) into four vertical zones (200–500 pc, 500–1000 pc, 1000–1500 pc, and 1500–2000 pc) as functions of time, while the yellow lines show the corresponding rate of change of cool gas mass in each zone. Bottom: same as top, but showing the cumulative net flux of cool gas since $t=15\,\rm Myr$ plus the cool gas mass at $t=15\,\rm Myr$ (blue), compared with the total cool gas mass (yellow).
@@ 4.2 Evolution of different phases in the wind
Figure 12 shows the distribution of the velocity along the minor axis, denoted as $V_{z}$, for the cool ($\rm{T}<2\times 10^{4}\,\rm{K}$, top panel), warm ($2\times 10^{4}$ to $2\times 10^{5}$ K, middle panel), and hot ($\rm{T}>2\times 10^{5}$ K, bottom panel) gas phases at different distances from the minor axis, r, in the FD simulation at 30 Myr. The solid and dashed orange lines denote the escape velocity calculated with and without DM halo, respectively. From left to right, the panels show results at heights of 500, 1000, 1500, and 2000 pc. A significant difference in the $V_{z}$-r space from z=500 to z=2000 pc is observed between the cool gas and the other two phases. At $z=500$ pc, most of the cool gas has a vertical velocity between $\sim 60\,\rm{km\cdot s^{-1}}$ and  $\sim\,300\rm{km\cdot s^{-1}}$, which increases moderately with height. This velocity is well below the escape velocity, suggesting that most of the cool gas in the wind will eventually fall back onto the disk.
[…]
@@
In contrast, the warm and hot gas in our simulations exhibits higher vertical velocities. Most warm gas has velocities between $200\sim 600\,\rm{km\cdot s^{-1}}$, while hot gas velocities range from $400\sim 1000\,\rm{km\cdot s^{-1}}$. A significant fraction of the warm and hot ionized gas in the wind can escape the galaxy’s potential well. However, the velocities of the hot gas in our simulations are still about $20\%-30\%$ lower than those inferred from X-ray observations of M82 (Strickland and Heckman 2009).
@@
We also analyzed the metallicity distribution of the three phases in the outflow. As shown by Figure 13, the three phases in the outflow exhibit similar metallicity distributions, with most gas having a metallicity of $Z\sim 0.5-2.0\,Z_{\odot}$. The hot gas at $z=500$ pc has a moderately higher metallcity. This level is approximately 50-100 times higher than the initial metallicity in the FD simulation. These features indicate that all gas in the outflow has undergone significant metal enrichment. As the second columns in Figure 6 and 7 show that, the ejecta of SNe have effectively enhanced the metallicity of cool, warm and hot phases in the central region of disc by $t=15.0$ Myr, when a significant portion of gas in disc began to flow outward to the wind.
To further confirm the origin of the cool gas in the outflow, we calculated the net mass flux of cool gas entering four height zones along the minor axis, 200-500 pc, 500-1000 pc, 1000-1500 pc and 1500-2000 pc, as a function of time. Figure 14 shows that the net mass flux of cool gas into the 500-1000 pc and 1000-1500 zones closely matches the growth rate of cool gas within them between $t=17.5$ and $t=30$ Myr. However, the change in cool gas mass slightly deviates from the net mass flux in the 1500-2000 pc zone, which could because of cooling of warm and hot gas. In the interface zone between disc and outflow ($200\,\rm{pc}\leq\,z<500\,\rm{pc}$ ), the mass change rate of cool gas also closely tracks the net mass flux between $t=20$ and $t=25$ Myr, but is higher at other times.
The lower panels of Figure 14 show that, since $t=15$ Myr (when the outflow begins to develop), the change in total cool gas mass between $z=500$ pc and $z=1500$ pc closely tracks the integrated net mass flux of cool gas. In contrast, the change in cool gas mass is slightly larger than the integrated net flux in the uppermost zone (z=1500z=1500–2000 pc) and moderately larger in the lower zone (z=200z=200–500 pc). This indicates that cooling from the warm and hot phases contributes little between $z=500$ and $z=1500$ pc, likely offset by ionization heating of the cool gas. By comparison, the contributions from in-situ cooling, which here refers to the condensation and precipitation of warm gas during its transport and mixing (Thompson et al., 2015), in the interface region (z=200z=200–500 pc) and in the uppermost zone are at the level of around $10$ to $30\%$ and around $10\%$, respectively.
Note that the cooling time of warm phase, fo gas with a number density of order $1\ cm^{-3}$ is around several to tens of thousands years, much shorter than the typical timescale of outflow. Therefore, both condensation and precipitation could occur efficiently in the warm phase. Conversely, processes such as stripping, heating, and ionization within the mixing layer can heat some of the cool gas to warm and hot phases. Together, these mechanisms can drive substantial mass exchange between the cool and warm phases; however, if the two processes approximately balance, the net change in the mass of each phase may remain small (Schneider et al. 2020; Abruzzo et al. 2021). In the present work, we did not employ passive scalars to track disk-origin cool gas, and therefore cannot directly quantify the instantaneous transfer rate between the cool and warm phases caused by heating, mixing, or cooling. Overall, our results suggest that the total mass of the cool phase in the outflow increases only modestly during transport. Any net transfer of material from hotter to cooler phases appears to be minor and may be largely offset by heating and ionization.
@@ 4.3 Overall feature of the outflow in various simulations
The launch of galactic winds varies significantly across different simulations. To investigate the impact of different feedback mechanisms, Figures 15 and 16 compare the face-on and edge-on views of the wind and disk at t = 25 Myr for simulations with different feedback configurations.
Compared to the FD simulation, the wind in the fSN simulation occupies a slightly larger volume and has a higher gas density, due to the greater total SN energy in fSN, which ejects more gas from the disc into the outflow. However, the intense SN feedback in the fSN simulation results in a more dispersed gas distribution in the central 1 kpc region. On the contrary, the nSN simulation fails to launch a galaxy-scale outflow. In addition, the gas disk morphology in the nSN simulation is dominated by prominent cool filamentary structures, resembling the fragmented state driven by gravitational instability described by Johnson and Gammie (2003). In contrast, disks in simulations with SN feedback are more turbulent, exhibiting well-developed density fluctuations across scales from tens of parsecs up to kiloparsecs. Although some smaller bubbles filled with warm gas, driven by stellar winds and radiation feedback, are present in the disc of nSN simulation, only a small amount of gas can escape into the halo, reaching heights of a few hundred to 1 kpc (Figure 16).
When both stellar wind and radiation feedback are further excluded (nFB simulation), the gas disk becomes even more compact. In this case, no gas from the disk can escape into the halo. Additionally, without SNe, the gas retains its initial metallicity. Figure 17 presents edge-on views of the outflow at 30 Myr for all simulations. Reducing the gas return fraction has a noticeable effect on the outflow morphology. When the fraction is decreased to 0.5, the outflow region becomes more extended, and further reduction to 0 leads to the appearance of low-density voids within the outflow. A higher initial gas metallicity results in a larger outflow volume due to enhanced star formation and SN feedback. Meanwhile, a less massive gas disk has a moderate impact on the morphology of the outflow.
@@
Figure 18: Left: the mass density function at $t=30$ Myr calculated along the minor axis in various simulations. Right: outflow rate at different heights along the minor axis. Each data point represents an average over a slit spanning the full box length and width, with a height of $\Delta z=150,\rm pc$.